OpenStudy (anonymous):

Solve the equation by completing the square: x2 + 8 = 6x Enter the smallest of the two zeros.

OpenStudy (nenadmatematika):

x^2-6x+8=0 (x-4)(x-2)=0 x=4 or x=2 so the answer is 2

OpenStudy (anonymous):